OpenStudy (anonymous):

Find x in this proportion. x/3=x=2/5 x=___? -1 1 2 3

OpenStudy (nenadmatematika):

write it again correctly :d

OpenStudy (anonymous):

oops sorry its x/3=x+2/5

OpenStudy (nenadmatematika):

something's wrong, you aren't given the correct choice....check again your writing

OpenStudy (anonymous):

nope this time is correct and also the answer choices

OpenStudy (nenadmatematika):

x=3x+6/5 , so 2x=-6/5, and x=-6/10=-3/5 so none of the above.....

OpenStudy (anonymous):

\[\frac{x}{3}=\frac{x+2}{5}\] \[5x=3(x+2)\] \[5x=3x+6\] \[2x=6\] \[x=3\]

OpenStudy (anonymous):

HMM ill take a guess, thanxx anyway=)

OpenStudy (anonymous):

i wrote it out, it is 3 no need to guess

OpenStudy (anonymous):

thanxx satellite73 =)

OpenStudy (nenadmatematika):

omg calistar I told you that didn't write it correctly...you must write (x+2)/5 if you want to get the correct answer....if you don't do that you'll make mistakes all the time....satellite gave the correct answer

OpenStudy (anonymous):

=( sorry..i didnt know that

OpenStudy (nenadmatematika):

it's ok :D